What is the acceleration of a 500 kg tricycle moving at 8 m/s takes a turn around a circle with a radius of 20m?


1
Expert's answer
2022-01-18T13:42:33-0500

"a=\frac{v^2}r=\frac{8^2}{20}=3.2~\frac m{s^2}."